OpenStudy (anonymous):

cancel each fraction to its simplest form 5 over 20 , 75 over 100, 15 over 18, 45 over 100,

OpenStudy (anonymous):

5/20=1/4 75/100=3/4 15/18=5/6 45/100=9/20